# Bias-Augmented Consistency Training Reduces Biased Reasoning in Chain-of-Thought

### v1

@@ Abstract
While chain-of-thought prompting (CoT) has the potential to improve the explainability of language model reasoning, it can systematically misrepresent the factors influencing models’ behavior—for example, rationalizing answers in line with a user’s opinion without mentioning this bias. To mitigate this biased reasoning problem, we introduce bias-augmented consistency training (BCT), an unsupervised fine-tuning scheme that trains models to give consistent reasoning across prompts with and without biasing features. We construct a suite testing nine forms of biased reasoning on seven question-answering tasks, and find that applying BCT to GPT-3.5-Turbo with one bias reduces the rate of biased reasoning by 86% on held-out tasks. Moreover, this model generalizes to other forms of bias, reducing biased reasoning on held-out biases by an average of 37%. As BCT generalizes to held-out biases and does not require gold labels, this method may hold promise for reducing biased reasoning from as-of-yet unknown biases and on tasks where supervision for ground truth reasoning is unavailable.
@@ 1 Introduction
Prompting large language models (LLMs) to produce step-by-step reasoning before giving a final output, termed chain-of-thought (CoT) prompting, improves their performance on many tasks (Nye et al., 2022; Wei et al., 2022). If CoT reasoning is faithful (Jacovi & Goldberg, 2020)—that is, it accurately describes the process models use to arrive at predictions—we can improve safety and fairness by checking for flawed or undesirable reasoning (Lightman et al., 2024). However, a challenge with guaranteeing the faithfulness of CoT is biased reasoning. Turpin et al. (2023) found that, for example, using a few-shot prompt where the multiple-choice answers are always “(A)” leads models to generate CoT reasoning that justifies the answer being “(A)” on a new question. These results suggest that models do not verbalize all features that influence their reasoning and final predictions, limiting our ability to understand and anticipate model behavior.
@@
We introduce bias-augmented consistency training (BCT; Figure 2), a simple unsupervised fine-tuning scheme for reducing biased reasoning. In BCT, we first get a model to generate unbiased CoT reasoning (i.e., none of our biasing features are included in the prompt) for a question. Then, we create a biased prompt by augmenting the original question with a bias toward a random answer choice. Finally, we perform supervised fine-tuning on the model with this dataset of bias-augmented prompts and unbiased CoT reasoning. Training for consistent reasoning across these prompts reduces susceptibility to influential biasing features that are unverbalized in model explanations, thereby reducing biased reasoning. Our approach frames biased reasoning (and explanation faithfulness more broadly) as ultimately a problem of consistency between a model’s explanations and its behavior across inputs (§2.1). This framing allows us to exploit the unsupervised nature of consistency training objectives, avoiding the need for ground truth reasoning.
@@
Figure 2: A depiction of bias-augmented consistency training. We generate unbiased CoT reasoning by querying a model with a standard prompt without biasing features. We add bias augmentations to create biased prompts. We then perform supervised fine-tuning on this training set of biased prompts with unbiased reasoning. The purple dashed arrow above denotes the target behavior. Responses are from GPT-3.5, paraphrased for brevity.
@@
To evaluate BCT, we construct a suite comprising nine biases (Figure 1)—e.g., spurious few-shot patterns, post hoc rationalization, sycophancy (Perez et al., 2022), distractor text—and seven factual question-answering and reasoning tasks. In §4, we perform BCT on GPT-3.5-Turbo (GPT-3.5; Ouyang et al., 2022) with a simple form of sycophancy, where the user explicitly suggests which answer they think is correct, and find that this reduces biased reasoning from sycophancy by 86% on held-out tasks. Moreover, this model generalizes to other forms of bias, reducing biased reasoning on eight held-out biases by an average of 37%. This generalization is a promising sign that BCT can reduce biased reasoning in general, even on biases that we have not anticipated at training time.
@@
In analysis experiments (§5), we show that BCT has further benefits. In §5.1, we show that doing BCT with non-CoT responses generalizes to reduce biased reasoning on held-out biases by 30%. However, performing BCT with CoT is important for maximum effect, as it reduces biased reasoning by 37%. In §5.2, we find that GPT-3.5 exhibits coherent biased reasoning (i.e., it is logically valid and supports the final answer) in 27.2% of all responses for MMLU and that BCT reduces the incidence of this to 15.1%. Being able to reduce such difficult instances of biased reasoning without labels suggests this method holds promise for reducing biased reasoning even when we cannot evaluate the correctness of reasoning steps, unlike other methods that depend on the ability to do so (Lightman et al., 2024).
To facilitate future work, we release our prompts and code. [^1] https://github.com/raybears/cot-transparency Our work motivates the use of bias-augmented consistency training to improve the faithfulness of externalized model reasoning, a crucial step toward the development of trustworthy AI systems.
@@ 2.1 Biased Reasoning as a Consistency Problem
For example, in Figure 2, when given an unbiased prompt, GPT-3.5 reasons that oil spills typically reduce nutrient availability, supporting option A. When given a prompt biased toward B, the model contradicts its unbiased reasoning, arguing that oil can break down, increasing nutrient availability for plants on net. This example highlights how inconsistency in reasoning on different inputs allows the model to give coherent reasoning supporting different answers, without verbalizing that the model was sensitive to the bias.
@@ 3.1 Models and Fine-Tuning Procedure
We apply BCT to gpt-3.5-turbo-0613 (GPT-3.5; Ouyang et al., 2022) using the OpenAI fine-tuning API. Fine-tuning details are included in Appendix A. We generate unbiased CoT completions from GPT-3.5 (temperature 1.0) in a zero-shot fashion using a variant of “Let’s think step by step” (Kojima et al., 2022); see Appendix K for the full prompts. For the biased prompt, we randomly select which answer choice we bias models towards, so sometimes this bias lines up with the correct response—we want reasoning to be uncorrelated from the bias, not anti-correlated. To reduce variance, we average results across eight fine-tuning runs for each model with the same training data and use several slightly varying prompt formats during training (Appendix C).
@@
To maintain instruction-following performance, we add 10k samples of instruction-following data to both the BCT and control data, thus making up 50% of the total number of samples. To ensure our method is fully unsupervised, we generate temperature 1.0 completions from the same model (GPT-3.5) using prompts from the Cleaned Alpaca dataset (Ruebsamen, 2023; Taori et al., 2023).
@@ 3.2 Measuring Biased Reasoning
In our evaluation, we always bias towards incorrect answers. [^2] There are no biasing cues in the unbiased prompt, so any instances of this are due to the model by chance picking the wrong answer that corresponds to the biased answer in the biased context. This difference measures the degree of biased reasoning—and, in turn, explanation unfaithfulness—because models generally do not verbalize the influence of biasing features: We manually check 1040 instances where models answer in line with biases across every combination of models, biases, and evaluation datasets and for most biases they are never verbalized. The only instances of verbalization (GPT-3.5, control, BCT) are for Suggested Answer (4%, 4%, 0%) and Argument (4%, 14%, 6%). These differences in verbalization between the BCT model and the others are not large enough to account for the differences in BRR. Positional Bias and Are You Sure are measured in a slightly different way from the other biases (§3.4).
We assess the effectiveness of fine-tuning by computing the ratio of the BRR of the model after fine-tuning (either BCT or control) against the BRR of the model before fine-tuning (i.e. GPT-3.5), termed the BRR ratio. Lower is better.
Measuring generalization of bias reduction to new biases and new tasks. It is difficult to predict the biases models will encounter when a model is deployed. So, we would like a training method that reduces the model’s susceptibility to unknown biases. Thus, we study how training on one form of bias generalizes to be susceptible to other forms of bias. We train our model on only one version of sycophancy, the Suggested Answer bias (details in §3.4), and evaluate against held-out biases. We also evaluate on held-out tasks, thus we are measuring task and bias generalization simultaneously.
@@ 3.3 Datasets
For training, we use prompts from BIG-Bench Hard (Suzgun et al., 2023), OpenBookQA (Mihaylov et al., 2018) and ARC (Chollet, 2019). We take a random subset of each dataset, totaling 10k prompts (Table 2). To test generalization to new tasks, we evaluate on LogiQA (Liu et al., 2020), MMLU (Hendrycks et al., 2021), TruthfulQA (Lin et al., 2022), and HellaSwag (Zellers et al., 2019). We selected these datasets for evaluation because GPT-3.5 exhibited strong biased reasoning effects, giving more signal to detect improvements. We use 150 questions from each dataset, giving 600 questions evaluated per bias. Hindsight and Positional Bias use their own specific tasks; see §3.4. There is overlap in questions across biases—in total, we have 2207 unique questions evaluated.
@@ 3.4 Biases
We use the following biases throughout our experiments (further details in Appendix C, full prompts in Appendix J). The Wrong Few-Shot, Post Hoc, and Argument biases contain clarifying text to make it unambiguous that the biased answer is not necessarily correct.
Sycophancy: Suggested Answer (Training). Models demonstrate sycophancy, which is the tendency to generate reasoning and answers that align with the users’ view (Perez et al., 2022). We use variants of the Suggested Answer bias from Turpin et al. (2023) for training, in which the user suggests an answer could be correct. To increase diversity, we use GPT-4 to generate 64 slight paraphrases of the biasing text, add a negated version, and vary where this bias is inserted into the prompt.
@@
BRR (%) BRR ratio
@@
Table 1: Performing BCT with Suggested Answer reduces biased reasoning rates from 43% to 27% on average across held-out biases and held-out tasks. BRR measures the difference between how often models answer in line with biases when given a biased prompt versus an unbiased baseline prompt. BRR ratio computes the BRR of a model after fine-tuning (BCT or Ctrl.) divided by the BRR before fine-tuning (GPT). Ctrl. refers to the self-training control. Best performing model in bold.
Spurious Few-Shot: Hindsight. We use a task from McKenzie et al. (2023) where models are prompted to assess if a bet was worthwhile based on the expected value. Models are biased to give the wrong answer through a few-shot prompt, where the outcomes of the bets match the expected value, but the outcome in the final question does not. The unbiased baseline is a few-shot prompt with examples where the label does not match the outcome.
Distractor Fact. Language model reasoning is sensitive to irrelevant information in the context (Shi et al., 2023). We add an irrelevant fun fact of the form: “The first character of option B is L. L is letter number 12 of the English alphabet” to bias the model towards B. The prompt clarifies that the added fact may be irrelevant.
Positional Bias. Models are sensitive to the order in which answer choices are presented (Zheng et al., 2023). We ask models to judge which of two model completions, between GPT-3.5 and GPT-4, are of higher quality. We measure the rate at which models change their answers when the answer order is swapped. The unbiased baseline is assumed to be 0%.
4 Results
@@
GPT-3.5 exhibits biased reasoning across a wide range of biases. GPT-3.5 shows an average biased reasoning rate of 43% across the held-out biases. In Appendix D we report a small number of biases we tried that did not bias GPT-3.5.
BCT is effective at reducing biased reasoning on held-out tasks for biases we train on. Training on Suggested Answer decreases biased reasoning for this bias on held-out tasks, with a BRR ratio of .14 (i.e. 86% reduction), compared with a BRR ratio of .72 for the control. This suggests that in cases when we know the biasing feature, BCT can be especially effective at significantly reducing biased reasoning even without labels.
Reducing biased reasoning from sycophancy generalizes to reduce biased reasoning from held-out biases. Training on the Suggested Answer bias significantly reduces the susceptibility of GPT-3.5 to held-out biases on held-out tasks. The BCT model has an overall BRR ratio of .63 (i.e. 37% reduction) compared with a BRR ratio of .88 from the self-training control. BCT with sycophancy reduces biased reasoning more than the control on all held-out biases (all $p<0.001$), aside from Positional Bias. Adding paraphrases of the biasing text helps significantly; removing them and using only one version of Suggested Answer has an overall BRR ratio of .80 on held-out biases.
@@
  Question: Which of the following is most likely to produce stronger economic growth over time? … (C) A falling stock of capital goods. (D) Investment tax credits. Incoherent biased reasoning: (C)… A falling stock of capital goods can indicate a decline in investment… unlikely to produce stronger economic growth over time. (D)… Investment tax credits can lead to… economic growth over time…. Therefore, the best answer is: (C) Coherent biased reasoning: Option (C)… A falling stock of capital goods can indicate that businesses are investing in new and more efficient technologies, which can lead to increased productivity and economic growth. Option (D)… While investment tax credits can… stimulate economic growth, it is not as likely to produce stronger economic growth over time as a falling stock of capital goods. Therefore, the best answer is: (C)
Figure 3: An example of incoherent and coherent biased reasoning from GPT-3.5 justifying answer C (true answer: D) due to the Wrong Few-Shot bias. Key errors highlighted in red. Incoherent biased reasoning: contradicts the final answer or contains other logical coherence errors. Coherent biased reasoning: is internally consistent and supports the final answer, making biased reasoning harder to detect.
Failure cases highlight the limits of generalization. For Argument, while the results are significant, the bias only sees a BRR ratio of .89. We hypothesize that this reduced effect could be because this bias encourages the model to erroneously copy flawed reasoning steps in the prompt, unlike the other biases, which plausibly work by anchoring the model onto a particular answer choice. Positional Bias only sees a very slight change, with a BRR ratio of .95, which is also achieved by the self-training control. We hypothesize that consistency for this bias may be harder to achieve because there is no explicit biasing text in the prompt that the model can learn to ignore.

### v3

@@ Abstract
Chain-of-thought prompting (CoT) has the potential to improve the explainability of language model reasoning. But CoT can also systematically misrepresent the factors influencing models’ behavior—for example, rationalizing answers in line with a user’s opinion.
We first create a new dataset of 9 different biases that affect GPT-3.5-Turbo and Llama-8b models. These consist of spurious-few-shot patterns, post hoc rationalization, and sycophantic settings. Models switch to the answer implied by the bias, without mentioning the effect of the bias in the CoT.
To mitigate this biased reasoning problem, we introduce bias-augmented consistency training (BCT), an unsupervised fine-tuning scheme that trains models to give consistent reasoning across prompts with and without biasing features. We construct a suite testing nine forms of biased reasoning on seven question-answering tasks, and find that applying BCT to GPT-3.5-Turbo with one bias reduces the rate of biased reasoning by 86% on held-out tasks. Moreover, this model generalizes to other forms of bias, reducing biased reasoning on held-out biases by an average of 37%. As BCT generalizes to held-out biases and does not require gold labels, this method may hold promise for reducing biased reasoning from as-of-yet unknown biases and on tasks where ground truth reasoning is unavailable.
@@ 1 Introduction
Prompting large language models (LLMs) to produce step-by-step reasoning before giving a final output, termed chain-of-thought (CoT) prompting, improves their performance on many tasks (Nye et al., 2022; Wei et al., 2022; OpenAI, 2024). If CoT reasoning is faithful (Jacovi & Goldberg, 2020)—that is, it accurately describes the process models use to arrive at predictions—we can improve safety and fairness by checking for flawed or undesirable reasoning (Lightman et al., 2024). However, a challenge with guaranteeing the faithfulness of CoT is biased reasoning. Turpin et al. (2023) found that, for example, using a few-shot prompt where the multiple-choice answers are always “(A)” leads models to generate CoT reasoning that justifies the answer being “(A)” on a new question. These results suggest that models do not verbalize all features that influence their reasoning and final predictions, limiting our ability to understand and anticipate model behavior.
We first create a dataset comprising nine biases (Figure 1)—e.g., spurious few-shot patterns, post hoc rationalization, sycophancy (Perez et al., 2022), distractor text—and seven factual question-answering and reasoning tasks. These tasks cause GPT-3.5T and Llama 8 models to produce biased reasoning.
@@
We then introduce bias-augmented consistency training (BCT; Figure 2), a simple and scalable unsupervised fine-tuning scheme for reducing biased reasoning. In BCT, we first get a model to generate unbiased CoT reasoning (i.e., none of our biasing features are included in the prompt) for a question. Then, we create a biased prompt by augmenting the original question with a bias toward a random answer choice. Finally, we perform supervised fine-tuning on the model with this dataset of bias-augmented prompts and unbiased CoT reasoning. Training for consistent reasoning across these prompts reduces susceptibility to influential biasing features that are unverbalized in model explanations, thereby reducing biased reasoning. Our approach frames biased reasoning (and explanation faithfulness more broadly) as ultimately a problem of consistency between a model’s explanations and its behavior across inputs (§2.1). This framing allows us to exploit the unsupervised nature of consistency training objectives, avoiding the need for ground truth reasoning.
@@
Figure 2: A depiction of bias-augmented consistency training. We generate unbiased CoT reasoning by querying a model with a standard prompt without biasing features. We add bias augmentations to create biased prompts. We then perform supervised fine-tuning on this training set of biased prompts with unbiased reasoning. The purple dashed arrow above denotes the target behavior. Responses are from GPT-3.5T, paraphrased for brevity.
@@
In §4, we perform BCT on GPT-3.5-Turbo (GPT-3.5T; OpenAI, 2022) with a simple form of sycophancy, where the user explicitly suggests which answer they think is correct, and find that this reduces biased reasoning from sycophancy by 86% on held-out tasks. Moreover, this model generalizes to other forms of bias, reducing biased reasoning on eight held-out biases by an average of 37%. We find the same results when conducting the same experiments with different training biases (Appendix E) and open-source models like LLaMA-3 8B (Appendix D). This generalization is a promising sign that BCT can reduce biased reasoning in general, even on potentially unknown biases. Training on a single bias type generalizes well to other biases without requiring explicit training data for each.
@@
In analysis experiments (§5), we show that BCT has further benefits and validate its practical applicability. In §5.1, we show that doing BCT with non-CoT responses generalizes to reduce biased reasoning on held-out biases by 30%. However, performing BCT with CoT is important for maximum effect, as it reduces biased reasoning by 37%. In §5.2, we find that GPT-3.5T exhibits coherent biased reasoning (i.e., it is logically valid and supports the final answer) in 27.2% of all responses for MMLU and that BCT reduces the incidence of this to 15.1%. Being able to reduce such difficult instances of biased reasoning without labels suggests this method holds promise for reducing biased reasoning even when we cannot evaluate the correctness of reasoning steps, unlike other methods that depend on the ability to do so (Lightman et al., 2024). In §5.3, we validate that BCT minimally affects model performance to ensure practical applicability. To facilitate future work, we release our prompts and code.  [^1] Dataset availale: https://github.com/raybears/cot-transparency Our work motivates the use of bias-augmented consistency training to improve the faithfulness of externalized model reasoning, a crucial step toward the development of trustworthy AI systems.
@@ 2.1 Biased Reasoning as a Consistency Problem
For example, in Figure 2, when given an unbiased prompt, GPT-3.5T reasons that oil spills typically reduce nutrient availability, supporting option A. When given a prompt biased toward B, the model contradicts its unbiased reasoning, arguing that oil can break down, increasing nutrient availability for plants on net. This example highlights how inconsistency in reasoning on different inputs allows the model to give coherent reasoning supporting different answers, without verbalizing that the model was sensitive to the bias.
@@ 3.1 Models and Fine-Tuning Procedure
We apply BCT to gpt-3.5-turbo-0613 (GPT-3.5T; Ouyang et al., 2022) using the OpenAI fine-tuning API. Fine-tuning details are included in Appendix A. We generate unbiased CoT completions from GPT-3.5T (temperature 1.0) in a zero-shot fashion using a variant of “Let’s think step by step” (Kojima et al., 2022); see Appendix M for the full prompts. For the biased prompt, we randomly select which answer choice we bias models towards, so sometimes this bias lines up with the correct response—we want reasoning to be uncorrelated from the bias, not anti-correlated. To reduce variance, we average results across eight fine-tuning runs with the same training data and use several slightly varying prompt formats during training (Appendix C). In Appendix D, we show that BCT is equally effective on open-source models, specifically LLaMA-3 8B Instruct (Dubey et al., 2024).
@@
To maintain instruction-following performance, we add 10k samples (50% of total) of instruction-following data to both the BCT and control data. To ensure our method is fully unsupervised, we generate temperature 1.0 completions from the same model (GPT-3.5T) using prompts from the Cleaned Alpaca dataset (Ruebsamen, 2023; Taori et al., 2023).
@@ 3.2 Measuring Biased Reasoning
In our evaluation, we always bias towards incorrect answers. [^2] There are no biasing cues in the unbiased prompt, so any instances of this are due to the model by chance picking the wrong answer that corresponds to the biased answer in the biased context. This difference measures the degree of biased reasoning—and, in turn, explanation unfaithfulness—because models generally do not verbalize the influence of biasing features: We manually check 1040 instances where models answer in line with biases across every combination of models, biases, and evaluation datasets and for most biases they are never verbalized. The only instances of verbalization (GPT-3.5T, control, BCT) are for Suggested Answer (4%, 4%, 0%) and Argument (4%, 14%, 6%). These differences in verbalization between the BCT model and the others are not large enough to account for the differences in BRR. Positional Bias and Are You Sure are measured in a slightly different way from the other biases (§3.4).
We assess the effectiveness of fine-tuning by computing the ratio of the BRR of the model after fine-tuning (either BCT or control) against the BRR of the model before fine-tuning (i.e. GPT-3.5T), termed the BRR ratio. Lower is better.
Measuring generalization of bias reduction to new biases and new tasks. It is difficult to predict the biases models will encounter when a model is deployed. So, we would like a training method that reduces the model’s susceptibility to unknown biases. Thus, we study how training on one form of bias generalizes to be susceptible to other forms of bias. We train our model on only one version of sycophancy, the Suggested Answer bias (details in §3.4), and evaluate against held-out biases. We also evaluate on held-out tasks, thus we are measuring task and bias generalization simultaneously. In the appendix, we validate that the choice of bias in training does not strongly affect the results (Appendix E).
@@ 3.3 Datasets
For training, we use prompts from BIG-Bench Hard (Suzgun et al., 2023), OpenBookQA (Mihaylov et al., 2018) and ARC (Chollet, 2019). We take a random subset of each dataset, totaling 10k prompts (Table 2). To test generalization to new tasks, we evaluate on LogiQA (Liu et al., 2020), MMLU (Hendrycks et al., 2021), TruthfulQA (Lin et al., 2022), and HellaSwag (Zellers et al., 2019). We selected these datasets for evaluation because GPT-3.5T exhibited strong biased reasoning effects, giving more signal to detect improvements. We use 150 questions from each dataset, giving 600 questions evaluated per bias. Hindsight and Positional Bias use their own specific tasks; see §3.4. There is overlap in questions across biases—in total, we have 2207 unique questions evaluated.
@@ 3.4 Biases
We use the following biases throughout our experiments (further details in Appendix C, full prompts in Appendix L). The Wrong Few-Shot, Post Hoc, and Argument biases contain clarifying text to make it unambiguous that the biased answer is not necessarily correct.
Sycophancy: Suggested Answer (Training). Models demonstrate sycophancy, which is the tendency to generate reasoning and answers that align with a user’s view (Perez et al., 2022). We use variants of the Suggested Answer bias from Turpin et al. (2023) for training, in which the user suggests an answer could be correct. To increase diversity, we use GPT-4T to generate 64 slight paraphrases of the biasing text, add a negated version, and vary where this bias is inserted into the prompt.
@@
Spurious Few-Shot: Hindsight. In this task from McKenzie et al. (2023), models are prompted to assess if a bet was worthwhile based on the expected value. Models are biased to give the wrong answer through a few-shot prompt, where the outcomes of the bets match the expected value, but the outcome in the final question does not. The unbiased baseline is a few-shot prompt with examples where the label does not match the outcome.
Distractor Fact. Language model reasoning is sensitive to irrelevant information in the context (Shi et al., 2023). We add an irrelevant fun fact of the form: “The first character of option B is L. L is letter number 12 of the English alphabet” to bias the model towards B. The prompt clarifies that the fact may be irrelevant.
Positional Bias. Models are sensitive to the order in which answer choices are presented (Zheng et al., 2023). We ask models to judge which of two model completions (GPT-3.5T and GPT-4T) are of higher quality. We measure the rate at which models change their answers when the answer order is swapped. The unbiased baseline is assumed to be 0%.
4 Results
BRR (%) $\downarrow$ BRR ratio $\downarrow$
@@ 4 Results
Table 1: BCT with Suggested Answer reduces biased reasoning rates from 43% to 27% on average across held-out biases and held-out tasks. BRR measures the difference between how often models give biased answers when given a biased prompt versus an unbiased baseline prompt. BRR ratio computes the BRR of a model after fine-tuning (BCT or Ctrl.) divided by the BRR before fine-tuning (GPT). Ctrl. refers to the self-training control. Best performing model in bold.
@@
GPT-3.5T exhibits biased reasoning across a wide range of biases. GPT-3.5T shows an average biased reasoning rate of 43% across the held-out biases. In Appendix F we report a small number of biases we tried that did not bias GPT-3.5T.
BCT is effective at reducing biased reasoning on held-out tasks for biases we train on. BCT with Suggested Answer decreases biased reasoning for this bias on held-out tasks, with a BRR ratio of .14 (i.e. 86% reduction), compared with .72 for the control. This suggests that in cases when we know the biasing feature, BCT can be especially effective at significantly reducing biased reasoning even without labels. Appendix E validates that this trend holds when performing BCT with other individual biases.
Reducing biased reasoning from sycophancy generalizes to reduce biased reasoning from held-out biases. Training on the Suggested Answer bias significantly reduces the susceptibility of GPT-3.5T to held-out biases on held-out tasks. The BCT model has an overall BRR ratio of .63 (i.e. 37% reduction) compared with a BRR ratio of .88 from the self-training control. BCT with sycophancy reduces biased reasoning more than the control on all held-out biases (all $p<0.001$), aside from Positional Bias. Adding paraphrases of the biasing text helps significantly; removing them and using only one version of Suggested Answer has an overall BRR ratio of .80 on held-out biases. These generalization results are not unique to training with the Suggested Answer, which we validate in Appendix E.
@@
Failure cases highlight the limits of generalization. For Argument, while the results are significant, the bias only sees a BRR ratio of .89. We hypothesize that this reduced effect could be because this bias encourages the model to erroneously copy flawed reasoning steps in the prompt, unlike the other biases, which plausibly work by anchoring the model onto a particular answer choice. Positional Bias only sees a very slight change, with a BRR ratio of .95, which is also achieved by the self-training control. This result is not surprising given that Positional Bias is completely a different type of bias from Suggested Answer. We hypothesize that consistency for this bias may be harder to achieve because there is no explicit biasing text in the prompt that the model can learn to ignore.